OpenStudy (anonymous):

what is the value of a1 for a geometic sequence with a4=40 and a6=160?

OpenStudy (tkhunny):

'r' is the common ratio. a is the first value in the sequence. a4 = a*(r^3) = 40 a6 = a*(r^5) = 160 Find that little guy, a. He's hiding in there.

OpenStudy (anonymous):

haha

OpenStudy (jmark):

The general geometric Progression is given as by making use of a4 and a6 value we find r value. \[\frac{a6 = ar^6-1}{a4 = ar^4-1}\] \[\frac{160 = ar^5}{40 = ar^3}\] \[\frac{160 = r^5}{40 = r^3}\] 4 = r5-3 4 = r2 take square root on both side √4 = √ r2 2 = r...........(1) use the value of r to find a. a4=a∗(2)4−1 40 = a23 40 = a 8 40/8 = a 5 = a.........(2) a1 = a∗(2)4−1 = 5 *2(1-1) = 5*20 =5*1 = 5 the value of a1 is 5. maths question paper for class 9 ICSE - http://icse.edurite.com/icse-question-papers/icse-maths-question-papers-for-class-9.html